### v3

@@ A.1.2 Hyperparameters
A.2 Prior Moments for Asymmetric Conjugate Prior
We set we $m_{\beta,i}=0$ to shrink the VAR coefficients to zero for data expressed either in first-differenced or growth rates, while for data in levels or log-levels, $m_{\beta,i}$ is set to be zero as well except the coefficient associated with the first own lag, which is set to be one. As for $V_{\beta,i}$, we take the following approach
$V_{\beta,i}=\left\{\begin{array}[]{ll}\frac{\kappa_{1}}{l^{2}s_{i}^{2}}&\text{for the coefficient on the }l\text{-th lag of variable }i\\ \frac{\kappa_{2}}{l^{2}s_{j}^{2}}&\text{for the coefficient on the }l\text{-thlag of variable }j~{}~{}(j\neq i)\\ \kappa_{3}&\text{for the intercept.}\end{array}\right.$ (A.5)
The hyperparameter $\kappa_{1}$ controls the overall shrinkage strength for coefficients on own lags, whereas $\kappa_{2}$ controls those on lags of other variables. Both will be selected by maximizing the associated marginal likelihood. We set $\kappa_{3}=100$, which implies essentially no shrinkage for the intercepts. [^25] Note that in this approach we are setting our priors on the VAR coefficients from the structural representation of the model. It would also be possible to set prior means and variances directly on the reduced-form coefficients, and then back-out the implied prior moments of the structural-form coefficients. Note this prior variance allows for cross-shrinkage, i.e. the parameter $\kappa_{2}$ can be different than $1$, which implies the degree of shrinkage can be different for lags of variables that are not the dependent variable of a given equation. This is therefore more general than the prior variance implied by the natural conjugate prior appearing in (A.1). The latter can be obtained as a special case by setting $\kappa_{2}=1$ (i.e. ruling out cross-shrinkage).
[…]
Appendix Appendix B Data and Transformations
Abbreviation Description Transf. code Medium Large X-large
PAYEMS All Employees: Total nonfarm 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
INDPRO IP Index 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
FEDFUNDS Effective Federal Funds Rate 1 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
UNRATE Civilian Unemployment Rate 1 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
RPI Real personal income 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
DPCERA3M086SBEA Real PCE 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
CMRMTSPLx Real Manu. and TradeIndustries Sales 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
CUMFNS Capacity Utilization: Manufacturing 1 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
CES0600000007 Avg Weekly Hours: Goods-Producing 1 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
HOUST Housing Starts, Total 4 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
S&P 500 S&P’s Common Stock Price Index: Composite 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
T1YFFM 1-Year Treasury C Minus FEDFUNDS 1 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
T10YFFM 10-Year Treasury C Minus FEDFUNDS 1 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
BAAFFM Moodys Baa Corporate Bond Minus FEDFUNDS 1 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
EXUSUKx U.S.-UK Foreign Exchange Rate 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
WPSFD49207 PPI: Final Demand: Finished Goods 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
PPICMM PPI: Metals and metal products 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
PCEPI Personal Consumption Expenditures 5 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
CES0600000008 Avg Hourly Earnings: Goods-Producing 6 $\mathrm{x}$ $\mathrm{x}$ $\mathrm{x}$
HWI Help-Wanted Index for U 1 $\mathrm{x}$ $\mathrm{x}$
HWIURATIO Help Wanted to Unemployed ratio 1 $\mathrm{x}$ $\mathrm{x}$
CLF16OV Civilian Labor Force 5 $\mathrm{x}$ $\mathrm{x}$
M1SL M1 Money Stock 5 $\mathrm{x}$ $\mathrm{x}$
M2SL M2 Money Stock 5 $\mathrm{x}$ $\mathrm{x}$
M2REAL Real M2 Money Stock 5 $\mathrm{x}$ $\mathrm{x}$
S&P div yield S&P’s Composite Common Stock: Dividend Yield 1 $\mathrm{x}$ $\mathrm{x}$
S&P PE ratio S&P’s Common Stock: Price/Earnings ratio 5 $\mathrm{x}$ $\mathrm{x}$
TB6MS 6-Month Treasury Bill 1 $\mathrm{x}$ $\mathrm{x}$
GS1 1-Year Treasury Rate 1 $\mathrm{x}$ $\mathrm{x}$
GS5 5-Year Treasury Rate 1 $\mathrm{x}$ $\mathrm{x}$
AAA Moody’s Seasoned Aaa Corporate Bond Yield 1 $\mathrm{x}$ $\mathrm{x}$
BAA Moody’s Seasoned Baa Corporate Bond Yield 1 $\mathrm{x}$ $\mathrm{x}$
OILPRICEx Crude Oil, spliced WTI and Cushing 5 $\mathrm{x}$ $\mathrm{x}$
INVEST Securities in Bank Credit at All Commercial Banks 5 $\mathrm{x}$ $\mathrm{x}$
@@ A.2 Prior Moments for Asymmetric Conjugate Prior
Notes: The dataset described in McCracken and Ng (2016) is available for download at https://research.stlouisfed.org/econ/. Column Transf. code indicates the transformation of a series $x_{t}$, where: (1) no transformation, (2) $\Delta x_{t}$, (5) $\Delta\log\left(x_{t}\right)$, (6) $\Delta^{2}\log\left(x_{t}\right)$ with $\Delta^{i}$ indicating $i$ th differences. Columns Medium, Large, and X-large refer to the different model sizes discussed in Section 3.

### v4

@@ A.1.2 Hyperparameters
Appendix Appendix B Are the Differences Statistically Significant?
In the main body of the paper, we have considered the distribution of RMSFEs relative to an AR benchmark where the distribution was computed across all of 120 series in the dataset. The fact that such distribution was to the left (right) of the threshold of 1 meant that overall a model was producing better (worse) forecasts than the AR model, but no assessment was made on whether each forecast individually was statistically different from the one produced by the benchmark. In this section we tackle this specific issue. To save space, we only consider a subset of 19 key macroeconomic variables. A full set of result, covering all 120 variables, is available upon request.
To provide a rough gauge of whether the RMSE ratios are significantly different from 1, we use the Diebold and Mariano (1995) t-statistic (DM). Our use of the DM test with models that are in some case nested is a deliberate choice. Monte Carlo evidence in Clark and McCracken (2012, 2015) indicates that, with nested models, the Diebold-Mariano test compared against normal critical values can be viewed as a somewhat conservative (conservative in the sense of tending to have size modestly below nominal size) test for equal accuracy in the finite sample. Differences that are statistically significant are denoted by one, two, or three stars, corresponding to significance levels of 10%, 5%, and 1%, respectively. The underlying p-values are based on t-statistics computed with a serial correlation-robust variance, using a rectangular kernel, h1h-1 lags, and the small-sample adjustment of Harvey et al. (1997).
Results for forecast horizons of one and three steps ahead are displayed in Table B.1. At these short forecast horizons the BVAR and factor model outperform the AR benchmark in many cases (12 and 14 cases respectively for BVAR and Factor model). In these instances, the forecasting gains are both large and statistically significant. In the remaining few cases in which the econometric models do not outperform the benchmark, the forecasting losses are small and never statistically significant. The TSLMs outperform the AR in fewer instances (5 for Moirai and 6 for TimesFM), also with gains that are both large and statistically significant. In the remaining cases, the forecasting losses are not statistically significant. As we have seen before, there is a tendency for TSLMs to produce the occasional more extreme results, with larger losses than those observed for the econometrics models.
$h=1$
BVAR Factor model NNAR BART Moirai Large TimesFM
PAYEMS 0.92 ** 0.87 ** 0.90 *** 0.97 0.83 *** 0.82 ***
INDPRO 0.92 *** 0.93 ** 0.92 *** 1.02 0.95 ** 0.94 **
FEDFUNDS 0.85 *** 0.91 2.34 1.34 1.00 1.23
UNRATE 0.92 *** 0.91 ** 1.03 0.94 1.04 1.11
RPI 0.98 0.98 1.02 1.05 1.00 1.04
DPCERA3M086SBEA 1.01 0.97 ** 1.04 1.03 1.02 1.06
CMRMTSPLx 1.01 0.97 1.05 1.01 1.01 1.04
CUMFNS 0.92 *** 0.96 1.93 0.96 1.09 1.48
CES0600000007 0.97 * 0.93 * 1.54 1.05 0.93 * 0.94 **
HOUST 0.99 0.94 *** 1.19 1.01 1.00 0.92 ***
S&P 500 1.03 1.00 1.05 1.03 1.03 1.02
T1YFFM 0.99 1.10 1.33 1.28 1.04 1.13
T10YFFM 0.94 ** 1.04 1.24 1.09 1.04 1.16
BAAFFM 0.93 * 0.97 1.19 1.07 1.01 1.28
EXUSUKx 1.04 1.01 1.06 1.05 1.02 1.01
WPSFD49207 0.98 1.03 1.00 1.03 1.01 1.01
PPICMM 1.06 0.99 1.06 1.02 1.03 1.04
PCEPI 1.04 0.98 0.98 1.02 0.99 0.95 *
CES0600000008 0.84 *** 0.78 *** 0.80 *** 0.91 *** 0.77 *** 0.78 ***
$h=3$
BVAR Factor model NNAR BART Moirai Large TimesFM
PAYEMS 0.89 * 0.75 *** 0.82 *** 0.99 0.77 *** 0.75 **
INDPRO 0.97 0.96 0.97 0.98 0.97 0.94
FEDFUNDS 0.87 *** 1.14 2.11 1.09 0.91 1.16
UNRATE 0.85 *** 0.78 * 0.99 0.98 0.91 0.78 **
RPI 0.99 * 0.99 * 0.99 1.02 1.00 1.01
DPCERA3M086SBEA 0.99 *** 0.99 0.98 ** 1.01 0.99 1.00
CMRMTSPLx 0.98 * 0.98 0.97 ** 1.01 1.00 1.01
CUMFNS 0.88 *** 0.91 1.89 1.07 0.93 1.32
CES0600000007 0.80 *** 0.78 *** 2.13 0.80 *** 0.79 *** 0.83 ***
HOUST 0.93 * 0.94 *** 1.48 1.04 1.01 0.78 ***
S&P 500 1.01 1.01 1.02 1.03 1.00 1.03
T1YFFM 1.00 1.18 1.14 1.32 1.03 1.12
T10YFFM 0.89 *** 1.08 1.09 1.05 0.96 1.13
BAAFFM 0.92 * 1.05 1.08 1.05 0.93 1.17
EXUSUKx 1.02 1.01 1.02 1.06 1.02 1.02
WPSFD49207 1.00 1.05 0.99 1.06 0.99 1.03
PPICMM 1.01 1.00 1.00 1.05 1.00 1.03
PCEPI 1.07 0.98 0.91 *** 0.98 0.89 *** 0.91 ***
CES0600000008 0.84 *** 0.76 *** 0.75 *** 0.91 *** 0.74 *** 0.75 ***
Table B.1: Forecast horizons 1 and 3 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is January 1985 to December 2019.
$h=6$
BVAR Factor model NNAR BART Moirai Large TimesFM
PAYEMS 1.00 0.99 ** 0.99 ** 0.81 ** 0.99 ** 0.99 *
INDPRO 1.00 1.00 0.99 1.01 0.99 0.98
FEDFUNDS 0.94 1.02 1.99 1.36 0.93 1.21
UNRATE 0.97 ** 0.91 * 1.01 0.85 0.95 0.94 **
RPI 1.00 1.00 1.00 1.03 1.00 1.01
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 1.00 1.00
CMRMTSPLx 1.00 1.00 1.01 0.98 1.01 1.02
CUMFNS 0.91 ** 0.97 1.69 1.01 0.92 1.14
CES0600000007 0.73 *** 0.78 *** 2.43 0.87 *** 0.76 *** 0.79 ***
HOUST 0.88 * 0.94 ** 1.80 0.99 0.98 0.66 **
S&P 500 1.01 1.01 1.01 1.05 1.00 1.03
T1YFFM 1.05 1.03 1.25 1.33 1.07 1.20
T10YFFM 0.89 ** 0.92 * 1.13 1.09 0.97 1.14
BAAFFM 0.94 1.00 1.08 1.07 0.93 1.15
EXUSUKx 1.01 1.00 1.03 1.09 1.00 1.02
WPSFD49207 1.01 1.03 0.98 1.07 0.99 1.02
PPICMM 1.02 1.01 1.01 1.04 1.01 1.06
PCEPI 1.03 0.93 0.89 *** 1.00 0.85 *** 0.85 ***
CES0600000008 0.85 *** 0.77 *** 0.76 *** 0.83 *** 0.76 *** 0.76 ***
$h=12$
BVAR Factor model NNAR BART Moirai Large TimesFM
PAYEMS 1.00 1.00 ** 1.00 1.00 1.00 0.99
INDPRO 1.00 1.00 1.00 1.04 1.01 0.99
FEDFUNDS 1.03 0.95 1.78 0.86 1.00 1.08
UNRATE 0.96 * 0.85 ** 1.03 0.99 0.96 0.89 *
RPI 1.00 1.00 1.00 1.01 1.00 1.00
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 1.00 1.00
CMRMTSPLx 1.00 1.01 1.01 1.03 1.02 1.02
CUMFNS 0.90 * 0.98 1.80 1.08 0.92 1.02
CES0600000007 0.73 *** 0.78 *** 3.08 0.80 *** 0.76 *** 0.78 ***
HOUST 0.86 0.99 2.20 1.18 0.96 0.48 **
S&P 500 1.00 1.00 1.01 1.04 1.01 1.05
T1YFFM 1.15 1.06 1.32 1.33 1.11 1.25
T10YFFM 0.97 0.87 *** 1.13 0.88 ** 1.03 1.15
BAAFFM 0.97 0.94 1.14 0.82 0.96 1.22
EXUSUKx 1.00 1.00 1.01 1.09 1.03 1.00
WPSFD49207 1.03 1.03 0.98 1.05 1.00 1.03
PPICMM 1.01 1.01 1.01 1.04 1.01 1.03
PCEPI 1.02 0.91 0.87 *** 0.94 0.82 *** 0.83 ***
CES0600000008 0.88 *** 0.80 *** 0.76 *** 0.91 ** 0.78 *** 0.80 ***
Table B.2: Forecast horizons 6 and 12 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is January 1985 to December 2019.
Results for forecast horizons of 6 and 12 step-ahead are displayed in Table B.2. At these horizons, for the 19 series in this medium dataset, improvements versus the benchmark are less clear-cut. In general, neither the econometric models nor the TSLMs systematically outperform the AR benchmark, even though it is worth noting that there are no instances in which they significantly under perform. Overall, Moirai is the model that more often produced better forecasts (in 13 cases at the 6-step ahead and 10 cases at the 12-step ahead) followed closely by the BVAR and factor model. For some series, there are significant gains. In particular the TSLMs produce good forecasts for Personal Consumption Expenditures (PCEPI) and Hourly Earnings (CES0600000008), with gains that are significant and larger than those obtained by the econometric models. The econometric models produce good forecasts of Unemployment rate (UNRATE) and Weekly hours (CES0600000007).
Appendix Appendix C Data and Transformations
Abbreviation Description Transf. code Small set Full set
PAYEMS All Employees: Total nonfarm 5 $\mathrm{x}$ $\mathrm{x}$
INDPRO IP Index 5 $\mathrm{x}$ $\mathrm{x}$
FEDFUNDS Effective Federal Funds Rate 1 $\mathrm{x}$ $\mathrm{x}$
UNRATE Civilian Unemployment Rate 1 $\mathrm{x}$ $\mathrm{x}$
RPI Real personal income 5 $\mathrm{x}$ $\mathrm{x}$
DPCERA3M086SBEA Real PCE 5 $\mathrm{x}$ $\mathrm{x}$
CMRMTSPLx Real Manu. and TradeIndustries Sales 5 $\mathrm{x}$ $\mathrm{x}$
CUMFNS Capacity Utilization: Manufacturing 1 $\mathrm{x}$ $\mathrm{x}$
CES0600000007 Avg Weekly Hours: Goods-Producing 1 $\mathrm{x}$ $\mathrm{x}$
HOUST Housing Starts, Total 4 $\mathrm{x}$ $\mathrm{x}$
S&P 500 S&P’s Common Stock Price Index: Composite 5 $\mathrm{x}$ $\mathrm{x}$
T1YFFM 1-Year Treasury C Minus FEDFUNDS 1 $\mathrm{x}$ $\mathrm{x}$
T10YFFM 10-Year Treasury C Minus FEDFUNDS 1 $\mathrm{x}$ $\mathrm{x}$
BAAFFM Moodys Baa Corporate Bond Minus FEDFUNDS 1 $\mathrm{x}$ $\mathrm{x}$
EXUSUKx U.S.-UK Foreign Exchange Rate 5 $\mathrm{x}$ $\mathrm{x}$
WPSFD49207 PPI: Final Demand: Finished Goods 5 $\mathrm{x}$ $\mathrm{x}$
PPICMM PPI: Metals and metal products 5 $\mathrm{x}$ $\mathrm{x}$
PCEPI Personal Consumption Expenditures 5 $\mathrm{x}$ $\mathrm{x}$
CES0600000008 Avg Hourly Earnings: Goods-Producing 6 $\mathrm{x}$ $\mathrm{x}$
HWI Help-Wanted Index for U 1 $\mathrm{x}$
HWIURATIO Help Wanted to Unemployed ratio 1 $\mathrm{x}$
CLF16OV Civilian Labor Force 5 $\mathrm{x}$
M1SL M1 Money Stock 5 $\mathrm{x}$
M2SL M2 Money Stock 5 $\mathrm{x}$
M2REAL Real M2 Money Stock 5 $\mathrm{x}$
S&P div yield S&P’s Composite Common Stock: Dividend Yield 1 $\mathrm{x}$
S&P PE ratio S&P’s Common Stock: Price/Earnings ratio 5 $\mathrm{x}$
TB6MS 6-Month Treasury Bill 1 $\mathrm{x}$
GS1 1-Year Treasury Rate 1 $\mathrm{x}$
GS5 5-Year Treasury Rate 1 $\mathrm{x}$
AAA Moody’s Seasoned Aaa Corporate Bond Yield 1 $\mathrm{x}$
BAA Moody’s Seasoned Baa Corporate Bond Yield 1 $\mathrm{x}$
OILPRICEx Crude Oil, spliced WTI and Cushing 5 $\mathrm{x}$
INVEST Securities in Bank Credit at All Commercial Banks 5 $\mathrm{x}$
@@
Notes: The dataset described in McCracken and Ng (2016) is available for download at https://research.stlouisfed.org/econ/. Column Transf. code indicates the transformation of a series $x_{t}$, where: (1) no transformation, (2) $\Delta x_{t}$, (5) $\Delta\log\left(x_{t}\right)$, (6) $\Delta^{2}\log\left(x_{t}\right)$ with $\Delta^{i}$ indicating $i$ th differences. Columns Medium and X-large refer to the different model sizes discussed in Section 3.